# Tapping into the Black Box: Uncovering Aligned Representations in Pretrained Neural Networks

### v1

@@ 5 Empirical validation
The goal of this section is to validate Hypothesis 1 empirically. In particular, if excitation pullbacks indeed approximate the decision boundary of a more regular linear model, then they should be better aligned with the data manifold, which should lead to better perceptual alignment. Moreover, we’d expect the excitation pullback to be highly robust to the choice of $\sigma$ across neurons, layers, and even architectures. This is because we’re interested in approximating the direction of $\gamma$ (not its exact value), and we assume $\gamma$ to be a binary tensor in $\mathbb{R}^{P_{1}}$, so it suffices to assign consistently higher scores to non-zero coordinates, implying robustness to hyperparameter choices.
To this end, we focus on popular and representative classes of ReLU-based architectures pretrained on ImageNet and available via the torchvision library, such that it is straightforward to replace most or all occurrences of ReLU and MaxPool2d (see Section 5.3) layers with our alternative variants, using a single recursive function over the child modules. Based on these criteria, we selected ResNet50, VGG11_BN, and DenseNet121.
Notably, determining the optimal choice of $\sigma$ for every neuron or layer in a given network is outside the scope of this work. Instead, we seek a single global set of hyperparameters that performs reasonably well across all layers of a given model - and ideally, across all selected architectures. Our goal is to evaluate whether the resulting excitation pullbacks indeed exhibit a substantially improved perceptual alignment compared to standard gradients.
@@ 5.1 Experimental setup
For each visualization, we construct a batch containing one random image from each of the five selected classes. We then compute excitation pullbacks for the classification neuron $y_{c}$ corresponding to each class, producing a $\times 5$ 5 grid: each row corresponds to an input image, and each column to the target class for which the pullback is computed. We also compute the pullback for a randomly selected ImageNet class (“ostrich”), shown in the last column.
We repeat the same setup, but instead of computing a single excitation pullback, we perform a rudimentary projected gradient ascent toward the logit (pre-activation) of each class, along the excitation pullback. Specifically, at each iteration we take a gradient step of $L_{2}$-norm 202020, whose direction is given by the excitation pullback. At each step we project the perturbed image on the ball of radius $\epsilon=100$ centered at the original image. We iterate this process 101010 times and plot both the final perturbed images and their differences from the originals. For contrast, we repeat the same procedure using vanilla input gradients in place of excitation pullbacks. Note that due to the choice of $\epsilon$, the perturbations are roughly fixed after 555 iterations (see Figure 1), the other 555 serving to further refine the perturbations and make them visually clearer.
@@ 5.2 Technical details
We can easily compute the excitation pullback $v_{\Gamma}$ by first computing the right pre-activations $z_{\ell}$ via $M_{G}$ (doing the ordinary forward pass) and then multiplying ReLU gradients by $\sigma(z_{\ell})$ in the backward pass, i.e. computing surrogate gradients for ReLU layers. This can be achieved in PyTorch (Paszke et al., 2019) by defining the appropriate torch.autograd.Function, see Listing 1. Alternatively, the more flexible Forward Gradient Injection can be used (Otte, 2024).
[…]
We download the pretrained weights via torchvision.models library. Specifically, we use resnet50, vgg11_bn and densenet121 models with the flag pretrained=True.222see Section 6.3 for more context. Gradients are visualised by torchvision.utils.make_grid(scale_each=True). We perform the gradient ascent with images rescaled to the [1,1][-1,1][ - 1 , 1 ] range (and then appropriately normalized before feeding to the model). We seed the dataloader with torch.Generator().manual_seed(314).
5.3 Excitation pullbacks via MaxPool2d layer
Most convolutional architectures use spatial non-linearities in the form of MaxPool2d layers which can be viewed as a generalization of ReLU, i.e. $\text{ReLU}(z)=\max(0,z)$ and both operations propagate gradients only through the local maxima. This motivates the application of surrogate gradients to these layers as well. Analogously to the excitation pullback modification for ReLU, we leave the forward pass of max pooling unchanged, but replace its gradient with that of a softmax pooling operation, using the strike-through trick. This indeed smoothens the pullbacks even further. Implementation details can be found in the Listing 2; in particular, we set the temperature parameter to 0.30.30.3, the same as in surrogate gradient for ReLU.
@@ 5.4 Plotting excitation pullbacks
We present the plots generated according to the procedure described in Section 5.1, i.e. Figures 23 and 4. As shown in the experiments, a single choice of the function $\sigma$ performs reasonably well across all considered architectures, indicating a substantial robustness of excitation pullbacks to hyperparameter selection. As noted in Section 5.2, the visualizations remain qualitatively similar across a wide range of $\sigma$ choices. Moreover, excitation pullbacks tend to highlight similar features across architectures, which suggests that the models learn comparable feature representations. Finally, the structure of the excitation pullbacks intuitively reflects the internal organization of each network, reinforcing our hypothesis that they indeed faithfully capture the underlying decision process of the model.
Figure 2: Top row: vanilla gradients; bottom row: excitation pullbacks. From left to right: ResNet50, VGG11_BN, DenseNet121.
Figure 3: Left: Image perturbations after 10-step projected gradient ascent along excitation pullbacks toward each of the classes (columns). Right: Difference between the perturbed and original images. From top to bottom: ResNet50, VGG11_BN, DenseNet121. One can clearly distinguish label-specific features highlighted by the model on every image. Compare with the same plots for vanilla gradients, Figure 4
@@ 6.1 Related work
Similar approach to ReLU networks from the path perspective has appeared in (Yadav et al., 2024) where authors study the so called Deep Linearly Gated Networks (DLGN) in which the gating function $G$ is computed by a parallel linear model $M$ with independent set of parameters $\theta$. The authors show that such models are considerably more interpretable than standard DNN’s and retain much of their performance. In their previous works (Lakshminarayanan and Singh, 2021) the authors define Neural Path Features (NPF) which are reminiscent of our gating functions (see Equation 3) and argue that “almost all the information learnt by a DNN with ReLU activations is stored in the gates”. They introduce Deep Gated Networks (DGN) and provide more evidence in their later work that gate adaptation is the key to generalisation (Lakshminarayanan and Singh, 2020), most notably, that “the winning lottery is in the gating pattern”, which reinforces our Hypothesis 1.
In (Linse et al., 2024) it is shown that setting the negative slope of LeakyReLU considerably higher in the backward pass than the forward pass leads to much better aligned global feature vizualisations via activation maximization, which can be directly explained by our results as LeakyReLU is equivalent to setting constant excitation function for each of the real half-lines (e.g. 1 for positive values and 0.3 for non-positive). In (Horuz et al., 2025) they generalize this idea to ReLU’s with different surrogate gradients, which is even closer to our approach, indicating that particular selection of surrogate ReLU gradient (e.g. B-SiLU) consistently improves generalization performance. Combined with our findings this leads to the open question, whether computing excitation pullbacks (instead of vanilla gradients) during training can improve generalization, and to what extent. Intuitively, one should not differentiate through the excitation gates explicitly as otherwise they may saturate and make the network overfit to the training data.
@@
It has been observed that networks trained to be robust to adversarial perturbations yield gradients that are often aligned with human perception, a phenomenon first observed in (Tsipras et al., 2019) and then referred to as perceptually-aligned gradients (PAG) in (Kaur et al., 2019). In (Srinivas et al., 2024) the authors attribute the PAG property to off-manifold robustness, which leads input gradients to lie approximately on the data manifold. This is resonates with our observation that excitation pullback $v_{\Gamma}$ provides a smoother approximation of the model’s decision boundary then the standard, hard-gated gradients (see Comments in Section 4) and therefore, by extension, may better align with the actual data manifold.
Our work suggests that even standard pretrained networks contain highly structured, semantically meaningful representations - but these get obfuscated by what we interpret as inherent gradient noise. This observation aligns with the broader perspective that explainability is not absent in modern networks, but merely hidden, and seemingly recoverable by principled methods such as ours.
@@ 6.2 On the possible biological plausibility of the implicit linear model
One may even venture a more speculative interpretation: the excitation pullbacks could be loosely associated with internal subjective representations, i.e. qualia - subjective, localized manifestations of experience that emerge from the brain activity taken as a whole - in a way similar to how excitation pullbacks emerge from the gating pattern in ReLU networks. This may open exciting direction in neuroscience and the philosophy of mind.
@@ 6.3 Limitations
Our experiments were run on pretrained torchvision models downloaded with the flag pretrained=True. This flag is now deprecated in the recent versions of the library, in particular it returns the older ResNet50_Weights.IMAGENET1K_V1. We’ve observed that it’s harder (but still possible) to achieve similar quality of excitation pullbacks for the more recent ResNet50_Weights.IMAGENET1K_V2 weights, at least when using a single $\sigma$ for all the neurons, which, as we’ve indicated earlier, might by an oversimplified approach. We speculate that this is because of different distribution of pre-activations in newer versions of the pretrained models.
The formal theory developed in this paper concerns specifically the ReLU networks. Intuitively, similar arguments should hold for other activation functions, in particular the the ones similar to ReLU, e.g. SiLU and GELU. Initial experiments for those architectures show promise but require certain refinement of the computation of $\Gamma$ to account for the more complex gradients. However, it’s not straightforward to apply our results to attention-based models.
@@ 7 Acknowledgments
This research was supported by a private grant. The reader might find it interesting to review the public weekly log from the research.

### v3

@@ 6 Empirical validation
This section describes how we visualize the excitation pullbacks. To this end, we focus on popular and representative classes of ReLU-based architectures pretrained on ImageNet and available via the torchvision library, such that it is straightforward to replace most or all occurrences of ReLU and MaxPool2d (see Section 6.3) layers with our alternative variants, using a single recursive function over the child modules. Based on these criteria, we selected ResNet50, VGG11_BN, and DenseNet121.
Notably, determining the optimal choice of $\sigma_{\ell}$ for every neuron or layer in a given network is outside the scope of this work. Instead, we seek a single global set of hyperparameters that performs reasonably well across all layers of a given model - and ideally, across all selected architectures. Our goal is to evaluate whether the resulting excitation pullbacks indeed exhibit a substantially improved perceptual alignment compared to standard gradients.
@@ 6.1 Experimental setup
For each visualization, we construct a batch containing one random image from each of the five selected classes. We then compute excitation pullbacks for the classification neuron $y_{c}$ corresponding to each class, producing a $5\times 5$ grid: each row corresponds to an input image, and each column to the target class for which the pullback is computed. We also compute the pullback for a randomly selected ImageNet class (“ostrich”), shown in the last column.
We repeat the same setup, but instead of computing a single excitation pullback, we perform a rudimentary projected gradient ascent toward the logit (pre-activation) of each class, along the excitation pullback. We perform the gradient ascent with images rescaled to the [1,1][-1,1] range (and then appropriately normalized before feeding to the model). Specifically, at each iteration we take a gradient step of $L_{2}$-norm $20$, whose direction is given by the excitation pullback. At each step we project the perturbed image on the ball of radius $\epsilon=100$ centered at the original image. We iterate this process $10$ times and plot both the final perturbed images and their differences from the originals. For contrast, we repeat the same procedure using vanilla input gradients in place of excitation pullbacks. Note that due to the choice of $\epsilon$, the perturbations are roughly fixed after $5$ iterations (see Figure 1), the other $5$ serving to further refine the perturbations and make them visually clearer.
@@ 6.2 Technical details
We can easily compute the excitation pullback $v_{\Gamma}$ by first computing the right pre-activations $z_{\ell}$ via $M_{G}$ (doing the ordinary forward pass) and then multiplying ReLU gradients by $\sigma_{\ell}(z_{\ell})$ in the backward pass, i.e. computing surrogate gradients for ReLU layers. This can be achieved in PyTorch (Paszke et al., 2019) by defining the appropriate torch.autograd.Function, see Listing 1. Alternatively, the more flexible Forward Gradient Injection can be used (Otte, 2024).
In the experiments we use a fixed global $\sigma$ for all the layers in all the networks, but ideally $\sigma_{\ell}$ could be adjusted for every hidden neuron separately, based on the neuron pre-activation distribution. We’ve observed that different choices of the global $\sigma$ produced similarly-looking pullback visualizations as long as the overall sigmoidal shape of the global $\sigma$ function remained the same. We’ve chosen $\sigma(z)=\operatorname{sigmoid(\frac{z}{temp})}$ for $\operatorname{temp}=0.3$ but any selection of the parameter temp from the approximate range $\text{temp}\in[0.15,0.5]$ seemed to work quite well across all the tested architectures, which implies the considerable robustness of the method to the hyperparameter choice. Notice that $\Gamma$ approaches hard gating $G$ as temp goes to zero and no gating at all as temp goes to infinity.
We use resnet50, vgg11_bn and densenet121 models with the flag pretrained=True from the torchvision.models library.555see Section 7.3 for more context. Gradients are visualised by torchvision.utils.make_grid(scale_each=True). We seed the dataloader with torch.Generator().manual_seed(314).
6.3 Excitation pullbacks through the MaxPool2d layer
Most convolutional architectures use spatial non-linearities in the form of MaxPool2d layers which can be viewed as a generalization of ReLU, i.e. $\text{ReLU}(z)=\max(0,z)$ and both operations propagate gradients only through the local maxima. This motivates the application of surrogate gradients to these layers as well. Analogously to the excitation pullback modification for ReLU, we leave the forward pass of max pooling unchanged, but replace its gradient with that of a softmax pooling operation, using the strike-through trick. This indeed smoothens the pullbacks even further. Implementation details can be found in the Listing 2; we set the temperature parameter to $0.3$, the same as in surrogate gradient for ReLU.
@@ 6.4 Plotting excitation pullbacks
We present the plots generated according to the procedure described in Section 6.1, i.e. Figures 23 and 4. As shown in the experiments, a single global choice of the function $\sigma$ performs reasonably well across all considered architectures, indicating a substantial robustness of excitation pullbacks to hyperparameter selection. As noted in Section 6.2, the visualizations remain qualitatively similar across a wide range of $\sigma$ choices. Moreover, excitation pullbacks tend to highlight similar features across architectures, which suggests that the models learn comparable feature representations. Finally, the structure of the excitation pullbacks intuitively reflects the internal organization of each network, reinforcing our hypothesis that they indeed faithfully capture the underlying decision process of the model.
Figure 2: Top row: vanilla gradients; bottom row: excitation pullbacks. From left to right: ResNet50, VGG11_BN, DenseNet121. Best viewed digitally.
@@ 7.1 Related work
Similar approach to ReLU networks from the path perspective has appeared in (Yadav et al., 2024) where authors study the so-called Deep Linearly Gated Networks (DLGN) in which the gating function $G$ is computed by a parallel linear model $M$ with independent set of parameters $\theta$. They show that such models are considerably more interpretable than standard DNN’s and retain much of their performance. We highlight the authors’ other relevant contributions in Section 5.1, motivating our Hypothesis 1. Overall, they arrive at similar formal constructions throughout their works but their main focus is on alternative architectures like DGNs; in contrast, we present a refined formal framework to study properties of standard architectures.
In (Linse et al., 2024) it is shown that setting the negative slope of LeakyReLU considerably higher in the backward pass than the forward pass leads to much better aligned global feature vizualisations via activation maximization, which can be directly explained by our results as LeakyReLU is equivalent to setting constant excitation function for each of the real half-lines (e.g. 1 for positive values and 0.3 for non-positive). In (Horuz et al., 2025) they generalize this idea to ReLU’s with different surrogate gradients, which is even closer to our approach, indicating that particular selection of surrogate ReLU gradient (e.g. B-SiLU) consistently improves generalization performance. Combined with our findings this leads to the open question, whether computing excitation pullbacks (instead of vanilla gradients) during training can improve generalization, and to what extent. Intuitively, one should not differentiate through the excitation gates explicitly as they may saturate and overfit; nevertheless, they can still be computed in the backward pass as alternative gradients.
@@
It has been observed that networks trained to be robust to adversarial perturbations yield gradients that are often aligned with human perception, a phenomenon first observed in (Tsipras et al., 2019) and then referred to as perceptually-aligned gradients (PAG) in (Kaur et al., 2019). In (Srinivas et al., 2024) the authors attribute the PAG property to off-manifold robustness, which leads input gradients to lie approximately on the data manifold. This resonates with our observation that excitation pullback $v_{\Gamma}$ provides a smoother approximation of the model’s decision boundary then the standard, hard-gated gradients and therefore, by extension, may better align with the actual data manifold.
@@ 7.2 On the possible biological plausibility of the implicit linear model
One may even venture a more speculative interpretation: the excitation pullbacks could be loosely associated with internal subjective representations, i.e. qualia - subjective, localized manifestations of experience that emerge from the brain activity taken as a whole - in a way similar to how excitation pullbacks emerge from the gating patterns in ReLU networks. It makes even more sense in the light of Section 5.4, where we interpret excitation pullbacks as the content of the network’s associative memory. This may open exciting direction in neuroscience and the philosophy of mind.
@@ 7.3 Limitations
Our experiments were run on pretrained torchvision models downloaded with the flag pretrained=True. This flag is now deprecated in the recent versions of the library, in particular it returns the older ResNet50_Weights.IMAGENET1K_V1. We’ve observed that it’s harder (but still possible) to achieve similar quality of excitation pullbacks for the more recent ResNet50_Weights.IMAGENET1K_V2 weights, at least when using a single global $\sigma$ for all the neurons, which, as we’ve indicated earlier, might by an oversimplified approach. We speculate that this is because of different distribution of pre-activations in newer versions of the pretrained models, which have been heavily regularized.
The formal theory developed in this paper concerns specifically the ReLU networks. Intuitively, similar arguments should hold for other activation functions, in particular the ones similar to ReLU, e.g. SiLU and GELU. Initial experiments for those architectures show promise but require certain refinement of the computation of $\Gamma$ to account for the more complex gradients. We haven’t yet explored the potential applicability of our method to attention-based models.
Additionally, we expect the convolutions to play rather significant role in the emergence of aligned excitation pullbacks, as spatial weight sharing seems to be crucial for the early path stabilization. This underlines the importance of an appropriate inductive bias for our results to carry over to other domains.
@@ 8 Acknowledgments
This research was conducted at the 314 Foundation [^6] https://314.foundation with support from a private grant awarded specifically for this study.